OpenStudy (anonymous):

how do i factor each polynomial? 18s-63

jimthompson5910 (jim_thompson5910):

what is the GCF of 18 and 63

OpenStudy (anonymous):

9

jimthompson5910 (jim_thompson5910):

how can you factor 18 where 9 is one factor

OpenStudy (anonymous):

9 goes into 18 twice...

jimthompson5910 (jim_thompson5910):

so 18 = 9*2

jimthompson5910 (jim_thompson5910):

which means 18s = 9*2s

jimthompson5910 (jim_thompson5910):

how can you factor 63 where 9 is one factor

OpenStudy (anonymous):

9^6?

jimthompson5910 (jim_thompson5910):

fill in the blank 9 times ______ = 63

OpenStudy (anonymous):

7

jimthompson5910 (jim_thompson5910):

so 63 = 9*7

jimthompson5910 (jim_thompson5910):

making 18s-63 turn into 9*2s - 9*7

jimthompson5910 (jim_thompson5910):

what's next

OpenStudy (anonymous):

is the answer 9^2-9^7?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

btw 9^2 is 9 squared, so you're using the wrong notation

OpenStudy (anonymous):

so what do i do next?

jimthompson5910 (jim_thompson5910):

hint: a*b - a*c = a*(b-c) this is the distributive property

OpenStudy (anonymous):

what do i do with the s?

jimthompson5910 (jim_thompson5910):

here's a similar example

jimthompson5910 (jim_thompson5910):

6x - 10 2*3x - 2*5 2(3x - 5)

OpenStudy (anonymous):

is 2 the gcf?

jimthompson5910 (jim_thompson5910):

that's my example

jimthompson5910 (jim_thompson5910):

sorry I have to go